# ExpliCa: Evaluating Explicit Causal Reasoning in Large Language Models

### v3

@@ Abstract
Martina Miliani1and Serena Auriemma1and Alessandro Bondielli1,2 Emmanuele Chersoni3  and Lucia C. Passaro2  and Irene Sucameli  and Alessandro Lenci1 1 CoLing Lab, Department of Philology, Literature, and Linguistics, University of Pisa, Italy 2 Department of Computer Science, University of Pisa, Italy 3 Department of Chinese and Bilingual Studies, The Hong Kong Polytechnic University martina.miliani@fileli.unipi.it, serena.auriemma@phd.unipi.it, alessandro.bondielli@unipi.it emmanuele.chersoni@polyu.edu.hk, lucia.passaro@unipi.it, irenesucameli@gmail.com, alessandro.lenci@unipi.it
@@ 1 Introduction
Understanding cause-effect relationships is one of the hallmarks of human cognition Pearl (2009). The question of whether Large Language Models (LLMs) truly comprehend causal relationships in natural language texts, or merely perform as ‘stochastic parrots’ Bender et al. (2021) by replicating statistical associations in their pretraining data, remains a topic of debate Zečević et al. (2023); Merrill et al. (2024). This question is crucial for the application of LLMs in domains that demand interpretive and inferential accuracy. The recent growth in causal research and benchmarking highlights the fundamental need for more reliable and interpretable models Chen et al. (2024). LLMs should be able to interpret not only cause-effect relations but also the relations between causal and temporal aspects (see Ning et al. (2018) as an example), which are often connected and overlapping (e.g., typically an effect temporally follows its cause).
@@
We conducted a nuanced evaluation of a group of LLMs on ExpliCa. Its goal is to shed light on several key aspects. First, we aim to estimate whether and to what extent LLMs can model and distinguish causal and temporal relations, similarly to humans. Second, we want to assess potential differences between LLMs’ competence and performance in our PCD task. Recent studies identified a discrepancy between the log-probability models assign to word sequences, which is referred to as their competence since it directly relates to their language modeling training objective, and models’ performance, which is given by the accuracy of their responses elicited via prompting, when tasked to provide an answer with instructions in natural language. This line of research suggests that prompting tends to underestimate the underlying linguistic knowledge of the models (Hu and Levy, 2023; Kauf et al., 2024). Therefore, we compare two metrics computed over connected sentences pairs: accuracy in prompting tasks, and perplexity. The latter is derived from the model’s log-probability and is used here as a proxy for the model’s acceptability decision, stemming from its underlying linguistic competence. Finally, we want to study these aspects across models of varying scales.
@@
Our primary contributions are shown in Fig. 1 and can be summarized as follows:
@@
  • we present a comprehensive evaluation of seven LLMs in total, comprising both commercial and open models (Sec. 5).  [^1] Data and code: github.com/Unipisa/explica
@@ 2 Related Work
The study of causality and its linguistic expressions garnered renewed and intensified interest, particularly in the context of evaluating the reasoning abilities of LLMs. Recent advancements led to the development of several specialized datasets, aimed at testing the causal reasoning of LLMs through hypothetical scenarios. Some studies evaluate LLMs on their ability to identify causal relations in structured data Cai et al. (2023). Notable works on unstructured data include CLadder Jin et al. (2023), which assesses causal reasoning using questions based on formal rules; CausalBench Wang (2024), used for tasks related to mathematics, coding, and textual data; and CausalNet Ashwani et al. (2024), which covers both causal and counterfactual questions. The focus of our work is on the extraction of causal events from unstructured textual data. Unlike in ExpliCa, the above mentioned datasets focus on implicit causal relations, which are not overtly expressed in linguistic structures.
@@
To evaluate LLMs, several datasets have also been annotated with explicit causal relationships between events in texts. These range from multilingual educational content, as in MECI (Lai et al., 2022), or financial news (Mariko et al., 2022), to more diverse sources, such as CREST Hosseini et al. (2021). Although these datasets do not explore the temporal dimensions of causality, many causal annotation schemas are derived from datasets that do annotate temporal relationships between events (e.g., BECauSE, Dunietz et al. 2017) or vice-versa (e.g., Temporal and Causal Reasoning, Ning et al. 2018), including those from news sources (e.g., Causal Time Bank, Mirza et al. 2014; Event StoryLine Corpus, Caselli and Vossen 2017), and from short commonsense narratives (e.g., CaTeRS, Mostafazadeh et al. 2016). However, such datasets do not leverage crowdsourcing annotation by native speakers for both causal and temporal relations as in ExpliCa. In our dataset, the ground truth is given by English native speakers’ annotation collected via crowdsourcing, with the aim of addressing the complexity of distinguishing truly causal from merely temporal relations between events. Furthermore, CLadder, CausalBench, and CausalNet are question-answering datasets that include scenarios involving causally related events. Another dataset of this kind is CRAB Romanou et al. (2023), which contains events taken from real-world situations together with their contextual information (i.e., news articles). Similarly, CausalProbe-2024 Chi et al. (2024) builds its scenarios from texts published after the release of the evaluated models, ensuring that the information was not available during model training. A key challenge in evaluating LLMs using datasets with direct textual annotations of causal relations is, in fact, the inherent ambiguity of their expression in natural language. For instance, linguistic markers such as and can signal either causality or temporality, depending on the context. This ambiguity can limit the effectiveness of such datasets in assessing causal reasoning.
@@
To overcome this limitation, ExpliCa has chosen a more controlled approach to causality evaluation by conducting a pairwise analysis of events, each expressed by a single sentence. The same strategy has been adopted in the COPA dataset Roemmele et al. (2011), where causality detection is framed as a task where the system must choose the most plausible alternative between two options. Similarly, e-CARE (Explainable Causal Reasoning, Du et al. 2022) includes over $21,000$ multiple-choice questions focused on causal reasoning, accompanied by conceptual explanations that clarify the underlying causal logic of each question. While these two datasets present instances of implicit causality, the BIG-Bench (Beyond the Imitation Game, Srivastava et al. 2022) initiative also models explicit causal reasoning. In this framework, the system must select the most plausible causal relationship between A because B and B because A. Similarly, in ExpliCa pairs of sentences from e-CARE and BIG-bench are joined in both directions using temporal and causal connectives. For example, the dataset contains both $A$ precedes/causes $B$ and $B$ precedes/causes $A$. This allows us to carefully analyze the models’ ability to discriminate between the related and yet very different relations of temporal precedence and causality. Furthermore, other linguistic cues, such as anaphoric references, have been removed. This design ensures that models are unlikely to rely on surface features, preventing the correct interpretation of causal markers from being inferred simply from the syntactic context.
@@
As in Kiciman et al. (2023) and Yu et al. (2025), many research works have focused on evaluating LLMs’ causal reasoning by using existing resources. In fact, some of the above datasets have become part of a broader evaluation framework called Causal evaluation of Language Models (CaLM, Chen et al. 2024). CaLM serves as a comprehensive benchmark for assessing LLMs’ causal reasoning capabilities. It comprises $126,334$ data samples and provides a foundational taxonomy of four modules: causal target, adaptation, metric, and error analysis. In relation to causal discovery, this framework addresses issues distinct from those targeted by ExpliCa, and it focuses solely on the analysis of LLM-generated responses. By contrast, in our work, we evaluated both the model outputs elicited via prompts, and the internal knowledge of LLMs, assessed through perplexity measurements. An overview of these datasets, along with a comparison with ExpliCa, is provided in Appendix A.
@@ 3 The ExpliCa Dataset
ExpliCa [^2] Release license details are in App. B. is designed to evaluate LLMs on commonsense causal reasoning through PCD tasks. It is composed of sentence pairs, where each sentence describes an event. Sentence pairs were in part adapted from sentences in existing datasets and in part manually crafted. Approximately a third of the sentence pairs were based on sentences from DeScript Wanzare et al. (2016), e-Care Du et al. (2022), and BIG-Bench Srivastava et al. (2022). [^3] Cf. Sec. 2 for a brief description of these datasets.
ExpliCa includes $600$ English sentence pairs, selected to be equally divided into three subsets: i.) causal subset, where the relationship is most likely causal (and possibly also temporal); ii.) temporal subset where the relation is expected to be only temporal; iii.) unrelated subset, including sentences that are thematically related but neither causally nor temporally. [^4] Note that the division was done by the authors during the creation of the dataset, but the final annotations provided as gold standard are based on the human ratings from the survey. Sentence pairs are linked through words that explicitly signal either a causal or temporal relationship. These connectives act as linguistic cues, enabling a causal or temporal interpretation of events based on the compositional meaning of the sentences. Other than the type of the relation between the events, the connectives specify the order of the events in the relations, which can be iconic — where the effect follows the cause and events are presented in their chronological order — or anti-iconic, where the order is reversed. We join each sentence pair with each connective, also considering the reverse order. In this way, we obtained $4,800$ unique items (600 pairs $\times$ 4 connectives $\times$ 2 orders).
An English native speaker examined a dataset sample to validate whether the connectives correctly express the nature of the relation and the order between the events. We detail the connectives below:
@@
Crucially, the connectives serve as the only linguistic cue for a causal or temporal interpretation of events. Other potential cues, such as causal verbs (e.g. cause, result, produce, affect etc.) were not included. Moreover, to avoid biasing humans’ and models’ decisions, anaphoric pronouns were not used. For example, instead of The coffee was bitter, The child spit it out, we overtly expressed the object in the second sentence: The coffee was bitter, The child spit the coffee out.
Then we applied a rigorous procedure to enrich, validate, and refine the dataset. First, we validated the dataset ensuring that the statistical association between words in sentence pairs does not significantly differ across the three subsets (i.e., Unrelated, Temporal, Causal). This validation was performed using Mutual Information (Sec. 3.1). Hence, we enriched each item with human acceptability ratings collected via crowdsourcing, and we set a threshold to the ratings for unrelated ones (Sec. 3.2). Finally, we checked for frequency biases in the dataset by analyzing the triplets {1st sentence verb, connective, 2nd sentence verb} in enTenTen Jakubíček et al. (2013) (Sec. 3.3).
@@ 3.1 Lexical Association Bias
If unrelated sentence pairs presented very different lexical elements compared to causally or temporally related ones, this might affect a LLM behavior, and lead to biased results. In order to address this aspect, we computed the statical association strength between pairs of lexemes (nouns, verbs, or adjectives), [^5] We used Stanza Qi et al. (2020) for PoS-tagging. one belonging to the first sentence and the other to the second sentence. To this aim, we used Mutual Information (MI) to quantify the strength of the relation between two linguistic units. More specifically, we adopted two variants: Pointwise Mutual Information (PMI) and Local Mutual Information (LMI) Church and Hanks (1990); Evert (2004). We averaged the MI scores of all the possible pairs of lexemes from a single sentence pair to obtain an item-level MI score, computed on the UkWaC corpus Ferraresi et al. (2008). We applied the Wilcoxon test to check if there were statistically significant differences in the item-level-MI scores across the Causal, Temporal, and Unrelated groups. We found that the statistical association of the sentence pairs in the three groups was not significantly different, both for LMI (W: $41,576$, p-value: $0.4312$) and PMI (W: $38,318$, p-value: $0.4009$). Thus, we can conclude that our dataset is free from lexical association biases.
@@ 3.2 Human Ratings
Each of the $4,800$ items in ExpliCa was annotated via crowdsourcing by 15 native English speakers. [^6] See App. C for details on the annotation procedure. We asked the participants to assess the acceptability of each item by giving a rating from 1 to 10. The ratings were then averaged for each item, obtaining the average acceptability rating for each connective in each sentence pair, in both event directions. We assigned to each sentence pair a relation type and a relation order label based on the connective deemed more acceptable for humans. For example, if then obtained the highest average rating for a sentence pair (in a specific direction), this is labeled as temporal type with an iconic order. Sentence pairs for which no connective had a rating higher than 6 and with mean rating below 5 are labeled unrelated. Tab. 1 summarizes the cardinality of the different classes in the dataset.
@@ 3.3 Frequency Bias
To control for such frequency biases in ExpliCa, we proceeded as follows. We used SpaCy,777spacy.io, English model en_core_web_sm. to lemmatize the sentences and perform part-of-speech (PoS) tagging, allowing us to identify the grammatical category of each linguistic unit. Then, we extracted the verb from each sentence, and the connective used to join them. For copulative verbs (e.g., is beautiful), we also considered the noun or the adjective following the copula. We retained particles in phrasal verbs (e.g., turn up), and in sentences containing multiple verbs (e.g., with embedded clauses), we considered up to two verbs, prioritizing those most salient to the sentence meaning (e.g., Michele chose the pizza he wanted to eat [want, eat]). Then, we used the SketchEngine APIs [^8] https://www.sketchengine.eu/apidoc/ to query the enTenTen21 corpus to compute the co-occurrence of the elements in the triplet {1st sentence verb, connective, 2nd sentence verb}. We divided the co-occurrences into frequency bins based on quartile ranges: Rare, Uncommon, Common, and Frequent. Fig. 2 shows the number of sentence pairs for each bin. While there are differences between classes, their distribution on the frequency spectrum shows no significant trends. [^9] Sentence count for each frequency bin in Tab. 4, App. D. This suggests that our dataset is relatively free from frequency biases.
@@ 4 Experiments
We evaluated models on ExpliCa using human ratings as ground truth across four tasks: three prompting tasks, and one perplexity evaluation. Prompting experiments were conducted under various settings, including few-shot (i.e., the prompt includes a few examples illustrating how to perform the task) and zero-shot (i.e., no examples are provided in the prompt) setups, and employing either greedy search [^10] More details on answer cleaning in App. E. or the Outlines framework Willard and Louf (2023) as decoding strategy for generating answers. Greedy search, or greedy decoding, selects the most probable next token at each step, while Outlines is a decoding framework that restricts model outputs to a predefined set of valid completions. This approach has been adopted in several recent studies and is now commonly used in the literature to control outputs of generative models Kauf et al. (2024). We used accuracy as evaluation metric. To study the effect of the model parameter scale, we then compared several models of the Qwen2.5 family on the acceptability rating task and perplexity. Finally, we compared the distribution of model ratings in the acceptability task and perplexity scores with human ratings, assessing their correlation. By analyzing the relationship between model ratings, perplexity, and human acceptability ratings, we aim to determine whether - and to what extent - LLMs can capture and distinguish causal and temporal relations in ways that align with human understanding.
@@
i.) acceptability ratings - we adopted the same design used in the survey with human participants (Sec. 3.2). Items for which the model failed to provide a rating were assigned a score of $-1$;
@@
We randomized the order of few-shot examples, options to choose from, and correct answers during inference. In the few-shot scenario, the models saw one example for each connective. [^12] Selected prompts and perplexity scores are in App. F.

### v4

@@ Abstract
Martina Miliani1and Serena Auriemma1and Alessandro Bondielli1,2 Emmanuele Chersoni3  and Lucia C. Passaro2  and Irene Sucameli  and Alessandro Lenci1 1 CoLing Lab, Department of Philology, Literature, and Linguistics, University of Pisa, Italy 2 Department of Informatics, University of Pisa, Italy 3 Department of Chinese and Bilingual Studies, The Hong Kong Polytechnic University martina.miliani@fileli.unipi.it, serena.auriemma@phd.unipi.it , alessandro.bondielli@unipi.it emmanuele.chersoni@polyu.edu.hk, lucia.passaro@unipi.it, irenesucameli@gmail.com , alessandro.lenci@unipi.it
@@
This is a preprint version of the paper published in Findings of ACL 2025.
@@ 1 Introduction
Understanding cause-effect relationships is one of the hallmarks of human cognition Pearl (2009). The question of whether Large Language Models (LLMs) truly comprehend causal relationships in natural language texts, or merely perform as ‘stochastic parrots’ Bender et al. (2021) by replicating statistical associations in their pretraining data, remains a topic of debate Zečević et al. (2023); Merrill et al. (2024). This question is crucial for the application of LLMs in domains that demand interpretive and inferential accuracy. The recent growth in causal research and benchmarking highlights the fundamental need for more reliable and interpretable models Chen et al. (2024). LLMs should be able to interpret not only cause-effect relations but also the relations between causal and temporal aspects (Ning et al. (2018) as an example), which are often connected and overlapping (e.g., typically an effect temporally follows its cause).
@@
We conducted a nuanced evaluation of a group of LLMs on ExpliCa. Its goal is to shed light on several key aspects. First, we aim to estimate whether and to what extent LLMs can model and distinguish causal and temporal relations, similarly to humans. Second, we want to assess potential differences between LLMs’ competence and performance in our PCD task. Recent studies (Hu and Levy, 2023; Kauf et al., 2024) identified a discrepancy between the linguistic competence of models, as measured by log-probability, and the accuracy of their responses elicited via prompting, with the latter method underestimating the models’ actual linguistic knowledge. Finally, we want to study these aspects across models of varying scales.
@@
The primary contributions of this paper are illustrated in Fig. 1 and can be summarized as follows:
@@
  • we present a comprehensive evaluation of seven LLMs in total, comprising both commercial and open models (Sec. 5). [^1] Data and code at: https://anonymous.4open.science/r/ExpliCa-6473/.
@@ 2 Related Work
The study of causality and its linguistic expressions garnered renewed and intensified interest, particularly in the context of evaluating the reasoning abilities of LLMs. Recent advancements led to the development of several specialized datasets, aimed at testing the causal reasoning of LLMs through hypothetical scenarios. Notable examples include CLadder Jin et al. (2023), which assesses causal reasoning using questions based on formal rules; CausalBench Wang (2024), used for tasks related to mathematics, coding, and textual data; and CausalNet Ashwani et al. (2024), which covers both causal and counterfactual questions. Unlike ExpliCa, these datasets focus on implicit notions of causality, which are not overtly expressed in the linguistic structures of the text.
@@
To evaluate LLMs, several datasets have also been annotated with explicit causal relationships between events in texts. These range from multilingual educational content, as in MECI (Lai et al., 2022), or financial news (Mariko et al., 2022), to more diverse sources, such as CREST Hosseini et al. (2021). Although these datasets do not explore the temporal dimensions of causality, many causal annotation schemas are derived from datasets that do annotate temporal relationships between events (e.g., BECauSE, Dunietz et al. 2017) or vice-versa (e.g., Temporal and Causal Reasoning, Ning et al. 2018), including those from news sources (e.g., Causal Time Bank, Mirza et al. 2014; Event StoryLine Corpus, Caselli and Vossen 2017), and from short commonsense narratives (e.g., CaTeRS, Mostafazadeh et al. 2016). However, such datasets do not leverage crowdsourcing annotation by native speakers for both causal and temporal relations as in ExpliCa. In our dataset, the ground truth is given by English native speakers’ annotation collected via crowdsourcing, with the aim of addressing the complexity of distinguishing truly causal from merely temporal relations between events. A key challenge in evaluating LLMs using datasets with direct textual annotations of causal relations is, in fact, the inherent ambiguity of their expression in natural language. For instance, linguistic markers such as and can signal either causality or temporality, depending on the context. This ambiguity can limit the effectiveness of such datasets in assessing causal reasoning.
@@
To overcome this limitation, ExpliCa has chosen a more controlled approach to causality evaluation by conducting a pairwise analysis of events, each expressed by a single sentence. The same strategy has been adopted in the COPA dataset Roemmele et al. (2011), where causality detection is framed as a task where the system must choose the most plausible alternative between two options. Similarly, e-CARE (Explainable Causal Reasoning, Du et al. 2022) includes over $21,000$ multiple-choice questions focused on causal reasoning, accompanied by conceptual explanations that clarify the underlying causal logic of each question. While these two datasets present instances of implicit causality, the BIG-Bench (Beyond the Imitation Game, Srivastava et al. 2022) initiative also models explicit causal reasoning. In this framework, the system must select the most plausible causal relationship between A because B and B because A. Similarly, in ExpliCa pairs of sentences from e-CARE and BIG-bench are joined in both directions ($A>B$; $B>A$), but using both temporal and causal connectives. This allows us to carefully analyze the models’ ability to discriminate between the related and yet very different relations of temporal precedence and causality. Furthermore, other linguistic cues, such as anaphoric references, have been removed. This design ensures that models are unlikely to rely on surface features, preventing the correct interpretation of causal markers from being inferred simply from the syntactic context.
@@
Finally, it is worth mentioning that some of the above datasets have become part of a broader evaluation framework called Causal evaluation of Language Models (CaLM, Chen et al. 2024). CaLM serves as a comprehensive benchmark for assessing LLMs’ causal reasoning capabilities. It comprises $126,334$ data samples and provides a foundational taxonomy of four modules: causal target, adaptation, metric, and error analysis. In relation to causal discovery, this framework addresses issues distinct from those targeted by ExpliCa, and it focuses solely on the analysis of LLM-generated responses. By contrast, in our work, we evaluated both the model outputs elicited via prompts, and the internal knowledge of LLMs, assessed through perplexity measurements.
@@ 3 The ExpliCa Dataset
ExpliCa [^2] Release license details are in App. A. is designed to evaluate LLMs on commonsense causal reasoning through PCD tasks. It is composed of sentence pairs, where each sentence describes an event. Sentence pairs were in part adapted from sentences in existing datasets and in part manually crafted. Approximately a third of the sentence pairs were based on sentences from DeScript Wanzare et al. (2016), e-Care Du et al. (2022), and BIG-Bench Srivastava et al. (2022). [^3] Cf. Sec. 2 for a brief description of these datasets.
ExpliCa includes $600$ English sentence pairs, selected to be equally divided into three subsets: i.) causal subset, where the relationship is most likely causal (and possibly also temporal); ii.) temporal subset where the relation is expected to be only temporal; iii.) unrelated subset, including sentences that are thematically related but neither causally nor temporally. [^4] Note that the division was done by the authors during the creation of the dataset, but the final annotations provided as gold standard are based on the human ratings from the survey. Sentence pairs are linked through words that explicitly signal either a causal or temporal relationship. These connectives act as linguistic cues, enabling a causal or temporal interpretation of events based on the compositional meaning of the sentences. Other than the type of the relation intercurring among the events, the connectives specify the order of the events in the relations, which can be iconic — where the effect follows the cause and events are presented in their chronological order — or anti-iconic, where the order is reversed. We join each sentence pair with each connective, also considering the reverse order. This way, we obtained $4,800$ unique items (600 pairs $\times$ 4 connectives $\times$ 2 orders).
An English native speaker examined a sample of the dataset to validate if the connectives correctly express the nature of the relation and the order between the events. We detail the connectives below:
@@
Crucially, the connectives serve as the only linguistic cue for a causal or temporal interpretation of events. Other potential cues, such as causal verbs (i.e., cause, result, produce, affect, etc.) were not included. Moreover, to avoid biasing humans’ and models’ decisions, anaphoric pronouns were not used. For example, instead of The coffee was bitter, The child spit it out, we overtly expressed the object in the second sentence: The coffee was bitter, The child spit the coffee out.
Then we applied a rigorous procedure to enrich, validate, and refine the dataset. First, we validated the dataset assuring that the statistical association between words in sentence pairs does not significantly differ across the three subsets (i.e., Unrelated, Temporal, Causal). This validation was performed using Mutual Information (Sec. 3.1). Hence, we enriched each item with human acceptability ratings collected via crowdsourcing, and we set a threshold to the ratings for unrelated ones (Sec. 3.2). Finally, we checked for frequency biases in the dataset by analyzing the triplets {1st sentence verb, connective, 2nd sentence verb} in enTenTen Jakubíček et al. (2013)(Sec. 3.3).
@@ 3.1 Lexical Association Bias
If unrelated sentence pairs presented very different lexical elements compared to causally or temporally related ones, this might affect a LLM behavior, and lead to biased results. In order to address this aspect, we computed the statical association strength via Pointwise Mutual Information (PMI) and Local Mutual Information (LMI) Church and Hanks (1990); Evert (2004) between pairs of lexemes (nouns, verbs, or adjectives), [^5] We used Stanza Qi et al. (2020) for PoS-tagging. one belonging to the first sentence and the other to the second sentence. MI scores were computed on UkWaC Ferraresi et al. (2008). We averaged the MI scores of all the possible pairs of lexemes from a single sentence pair to obtain an item-level MI score. We applied the Wilcoxon test to check if there were statistically significant differences in the item-level-MI scores across the Causal, Temporal, and Unrelated groups. We found that the statistical association of the sentence pairs in the three groups was not significantly different, both for LMI (W: $41,576$, p-value: $0.4312$) and PMI (W: $38,318$, p-value: $0.4009$). Thus, we can conclude that our dataset is free from lexical association biases.
@@ 3.2 Human Ratings
Each of the $4,800$ items in ExpliCa was annotated via crowdsourcing by 15 native English speakers. [^6] See App. B for details on the annotation procedure. We asked the participants to assess the acceptability of each item by giving a rating from 1 to 10. The ratings were then averaged for each item, obtaining the average acceptability rating for each connective in each sentence pair, in both event directions. We assigned to each sentence pair a relation type and a relation order label based on the connective deemed more acceptable for humans. For example, if then obtained the highest average rating for a sentence pair (in a specific direction), this is labeled as temporal type with an iconic order. Sentence pairs for which no connective had a rating higher than 6 and with mean rating below 5 are labeled unrelated. Tab. 1 summarizes the cardinality of the different classes in the dataset.
@@ 3.3 Frequency Bias
To control for such frequency biases in ExpliCa, we proceeded as follows. We PoS-tagged and lemmatized the sentence pairs with SpaCy.777spacy.io, English model en_core_web_sm. Then, we extracted the verb from each sentence, and the connective used to join them. For copulative verbs, we also considered the noun or the adjective following the copula. We retained particles in phrasal verbs (e.g., turn up), and in sentences containing multiple verbs (e.g., with embedded clauses), we considered up to two verbs, prioritizing those most salient to the sentence meaning (e.g., Michele chose the pizza he wanted to eat [want, eat]). Then, we used the SketchEngine APIs [^8] https://www.sketchengine.eu/apidoc/ to query the enTenTen21 corpus to compute the co-occurrence of the elements in the triplet {1st sentence verb, connective, 2nd sentence verb}. We divided the co-occurrences into frequency bins based on quartile ranges: Rare, Uncommon, Common, and Frequent. Fig. 2 shows the number of sentence pairs for each bin. While there are differences between classes, their distribution on the frequency spectrum shows no significant trends. [^9] Sentence count for each frequency bin in Tab. 3, App. C. This suggests that our dataset is relatively free from frequency biases.
@@ 4 Experiments
We evaluated models on ExpliCa using human ratings as ground truth across four tasks: three prompting tasks, and one perplexity evaluation. Prompting experiments were conducted under various settings, including few-shot and zero-shot setups, and employing either greedy search [^10] More details on answer cleaning in App. D. or the Outlines framework Willard and Louf (2023) for generating answers. We used accuracy as evaluation metric. To study the effect of the model parameter scale, we then compared several models of the Qwen2.5 family on the acceptability rating task and perplexity. Finally, we compared the models’ rating distributions in the acceptability tasks to the human ones, and assessed their correlation.
@@
i.) acceptability ratings - we adopted the same design used in the survey with human participants (Sec. 3.2). Items for which the model failed to provide a rating were assigned a score of 1-1;
@@
We randomized the order of few-shot examples, options to choose from, and correct answers during inference. In the few-shot scenario, the models saw one example for each connective. [^12] Selected prompts and perplexity scores are in App. E.